A man uses a 900watt electric iron to press his cloth for an average of 4 hrs a week for 5 weeks. If the cost of electrical energy is 30k per unit(kWh), calculate the cost of energy consumed
1
Expert's answer
2020-06-25T10:11:31-0400

The total amount of energy can be calculated as the product of the power and the time the iron worked.

So the energy will be "900\,\mathrm{W}\cdot 4\,\mathrm{h}\cdot5 = 18000\,\mathrm{W\cdot h} = 18\,\mathrm{kW h}" .

Therefore, the cost will be "18\,\mathrm{kW h} \cdot 30\,\mathrm{k/kWh} = 540\,\mathrm{k}."